OpenStudy (anonymous):

Calculate the time required to send a 56 byte packet from Sydney to Melbourne if the data rate is155mbps, the distance traveled is 1000km and the speed of propagation is 2.8*10^8m/s.

OpenStudy (anonymous):

Let's say the packet starts being transmitted at t=0. The propagation time is d/c = 1000e3/2.8e8 = 3.5714e-3 seconds. This is when the leading edge of the packet just starts arriving at the receiver. The duration of the packet is 56 / 115e6 = 4.87e-7 seconds. The entire packet will have arrived at the receiver at the sum of these two times, or 3.5719 msec.